# Understanding Oversmoothing in GNNs as Consensus in Opinion Dynamics

### v1

@@ Abstract
In contrast to classes of neural networks where the learned representations become increasingly expressive with network depth, the learned representations in graph neural networks (GNNs), tend to become increasingly similar. This phenomena, known as oversmoothing, is characterized by learned representations that cannot be reliably differentiated leading to reduced predictive performance. In this paper, we propose an analogy between oversmoothing in GNNs and consensus or agreement in opinion dynamics. Through this analogy, we show that the message passing structure of recent continuous-depth GNNs is equivalent to a special case of opinion dynamics (i.e., linear consensus models) which has been theoretically proven to converge to consensus (i.e., oversmoothing) for all inputs. Using the understanding developed through this analogy, we design a new continuous-depth GNN model based on nonlinear opinion dynamics and prove that our model, which we call behavior-inspired message passing neural network (BIMP) circumvents oversmoothing for general inputs. Through extensive experiments, we show that BIMP is robust to oversmoothing and adversarial attack, and consistently outperforms competitive baselines on numerous benchmarks.
Machine Learning, ICML
@@ 1 Introduction
Graphs can be used to describe observations in diverse contexts, such as molecular dynamics (Fang et al., 2022), biological behavior, and computer networking. Graph neural networks (GNNs) leverage this structure to learn the features of the nodes and edges within these systems. Numerous GNN architectures have been proposed, including graph convolutional networks, (GCN) (Kipf & Welling, 2016), graph attention networks (GAT) (Veličković et al., 2017) GraphSAGE (Hamilton et al., 2017), and message passing neural networks (Gilmer et al., 2017). These methods respect the permutation equivariance of the underlying graph through constraints on the node and edge aggregation functions. However, it is exactly the structure of these aggregation functions that leads to oversmoothing, a phenomenon where network features become increasingly indistinguishable with increasing network depth (Oono & Suzuki, 2019; Nt & Maehara, 2019).
Recent works address oversmoothing in three different ways (Rusch et al., 2023): using residual connections (Li et al., 2019; Chen et al., 2020b; Xu et al., 2018; Liu et al., 2020); regularizing node features (Li et al., 2019; Chen et al., 2020b; Xu et al., 2018; Liu et al., 2020); and using physical inductive biases as message-passing functions.
This last approach is commonly used in continuous-depth GNN models where node features are propagated forward in time using a continuous-time differential equation rather than the discrete layerwise transformations of conventional message-passing architectures. Dynamical equations such as the diffusion equation (Chamberlain et al., 2021b; Thorpe et al., 2022), wave equation (Eliasof et al., 2021), coupled oscillator equations (Rusch et al., 2022; Nguyen et al., 2024), Allen-Cahn equation for particle systems (Wang et al., 2022), and geometrical equations like Beltrami flows (Chamberlain et al., 2021a) and Racci flows (Topping et al., 2021) have been successfully applied in this context.
Instead of using physical inductive biases to model the evolution of node features, we propose the use of opinion dynamics, a multi-agent systems model that captures the evolution of agent opinions (or node features in our case) over time (Zha et al., 2020; Dong et al., 2021; Montes de Oca et al., 2010). While the evolution of agent opinions converge for all initial conditions under a linear opinion dynamics model, the opinions of agents can have diverse outcomes under the nonlinear opinion dynamics (NOD) model (Bizyaeva et al., 2022; Leonard et al., 2024). This flexibility is due to the nonlinearity in NOD which introduces a bifurcation, allowing opinions to move toward consensus or dissensus (i.e., not consensus) depending on the initial conditions and network connectivity.
In this paper, we propose an analogy between oversmoothing in GNNs and opinion consensus in opinion dynamics models leading to the following contributions:
  • We prove that the dynamics of recent continuous-depth GNN models are equivalent to linear opinion dynamics and will consequently exhibit oversmoothing for all initial conditions,
  • We design a novel continuous-depth GNN model using a nonlinear opinion dynamics inductive bias, and prove that our behavior-inspired message passing (BIMP) model has greater expressive capacity than other continuous-depth GNN models,
  • We perform extensive empirical analysis and show that our BIMP model consistently outperforms competitive baseline models.
@@ 2 Related Work
Oversmoothing in graph neural networks.
Oversmoothing problem is widely noticed and discussed in (Li et al., 2018; Oono & Suzuki, 2019; Nt & Maehara, 2019). As the learned embeddings get similar, oversmoothing can be interpreted as an expressivity decay which is often analyzed using the Weisfeiler-Lehman test (Huang & Villar, 2021). Recent works provide precise definitions of oversmoothing through various perspectives, including the feature distance (Chen et al., 2020a), Dirichlet energy (Cai & Wang, 2020; Zhou et al., 2021; Rusch et al., 2022, 2023) and information gain (Zhou et al., 2020). Given these definitions, however, finding an efficient method to mitigate oversmoothing remains challenging and attracts significant attention in GNNs research.
Continuous-depth graph neural networks.
Continuous-depth networks such as NODE (Chen et al., 2018) define the network depth implicitly via the simulation of the given differential equations. Further work presented a framework (Poli et al., 2019) where inputs propagate through a continuum of GNN layers for a continuous-depth GNN. Extensions to this concept embeds many physical systems as the differential equation in continuous-depth GNNs. Rusch et al. (2022) leverages the stability characteristics in coupled oscillators to tackle the oversmoothing problem; Wang et al. (2022) incorporates the Allen-Cahn reaction-diffusion to achieve stricter lower-bounds on the Dirichlet energy for GNNs; Nguyen et al. (2024) models the underlying dynamics as oscillators in a Kuramoto system to reformulate oversmoothing as a problem of synchronization.
Opinion dynamics.
Since opinions can be interpreted as latent preferences, modeling opinion dynamics provides insights into understanding and predicting agents’ behavior. Specifically, consensus dynamics (Bullo, 2018) is commonly used to model the coordination of multi-agent systems. Leonard et al. (2010) describe a linear consensus dynamics for coordinated surveying with underwater gliders. Justh & Krishnaprasad (2005) use the consensus dynamics framework to formulate rectilinear and circular control laws for a multi-vehicles system. In (Leonard et al., 2007), the authors use consensus dynamics to improve data collection coverage in a mobile sensor network. Ballerini et al. (2008) use consensus dynamics to understand the robustness of starling flocking behavior. However, there are several draw backs to using a linear model for opinion formulation; specifically, the naive implementation results in dynamics that only converge to consensus (Altafini, 2012; Dandekar et al., 2013).
Nonlinear opinion dynamics.
@@
The noted short comings of consensus dynamics models are resolved in the nonlinear opinion dynamics (Leonard et al., 2024). The nonlinearity of the model results in bifurcations (Golubitsky et al., 2012; Guckenheimer & Holmes, 2013; Strogatz, 2018) and opinions can evolve to dissensus rapidly even under weak input signals (Bizyaeva et al., 2022). Nonlinear opinion dynamics have been shown to able to model a variety of systems such as political polarization (Leonard et al., 2021); group decision making (Dong et al., 2021); multi-agent control (Montes de Oca et al., 2010; Leonard et al., 2010; Cathcart et al., 2023); and relational inference (Yang et al., 2024). Our proposed network seeks to integrate nonlinear opinion dynamics as the behavior-inspired inductive bias in a continuous-depth GNN.
3 Preliminaries
In this section we describe our notation and review nonlinear opinion dynamics (Leonard et al., 2024).
3.1 Notation
Throughout this text we use lower-case letters for scalar valued variables, boldface lower-case letters for vectors, and boldface upper-case letters for matrices. We define graphs as the set $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n=|\mathcal{V}|$ nodes and $m=|\mathcal{E}|$. For multi-agent system evolving under opinion dynamics models we use $N_{\mathrm{a}}$ to denote the number of agents and $N_{\mathrm{o}}$ to denote the number of options.
3.2 Nonlinear opinion dynamics
Nonlinear opinion dynamics captures the evolution of agent opinions in a multi-agent system. Opinions for each option can be positive, neutral, or negative, where the magnitude correlates to the opinion strength. Changes to these opinions are affected by extrinsic and intrinsic parameters. Specifically, agent $i$’s opinion about option $j$ can be described by the following nonlinear differential equation proposed in (Bizyaeva et al., 2022; Leonard et al., 2024),
$\dot{x}_{ij}=-d_{ij}x_{ij}+S\Bigg{(}u_{i}\Bigg{(}\alpha^{j}_{i}x_{ij}+\sum^{N_{\mathrm{a}}}_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}a^{\mathrm{a}}_{ik}x_{kj}+\\ \sum^{N_{\mathrm{o}}}_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}a^{\mathrm{o}}_{jl}x_{il}+\sum^{N_{\mathrm{a}}}_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}\sum^{N_{\mathrm{o}}}_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{jl}x_{kj}\Bigg{)}\Bigg{)}+b_{ij}.$ (1)
The parameters $d_{ij}\geq 0$, $u_{i}>0$, and $\alpha_{i}^{j}\geq 0$ are intrinsic to each agent. The damping term $d_{ij}$ describes the resilience of agent $i$ to forming an opinion on option $j$; the attention $u_{i}$ represents the attentiveness of agent $i$ to the opinions of other agents; and the self-reinforcement $\alpha_{i}^{j}$ defines the resistance of agent $i$ to changing the opinion for option $j$.
The parameters $a^{\mathrm{a}}_{jk}$, $a^{\mathrm{o}}_{jl}$, and $b_{ij}$, and saturating function $S\left(\cdot\right)$ are extrinsic to the agents. The environmental $b_{ij}$ represents the impact of the environment on the agent’s opinion for option $j$. The saturating function $\mathcal{S}$ is selected so that $S\left(0\right)=0$, $S^{\prime}\left(0\right)=1$, and $S^{\prime\prime\prime}\left(0\right)\neq 0$. We use a hyperbolic tangent for the saturation function in our method.
The communication network graph $\mathcal{G}^{\mathrm{a}}=(\mathcal{V}^{\mathrm{a}},\mathcal{E}^{\mathrm{a}},\mathbf{A}^{\mathrm{a}})$ describes how agents are able to exchange their opinions over time. The communication adjacency matrix $\mathbf{A}^{\mathrm{a}}=\left[a^{\mathrm{a}}_{ik}\right]\in\mathbb{R}^{N_{\mathrm{a}}\times N_{\mathrm{a}}}$ defines the communication strength between agent $i$ and agent $k$. Similarly, the belief system graph $\mathcal{G}^{\mathrm{o}}=\left(\mathcal{V}^{\mathrm{o}},\mathcal{E}^{\mathrm{o}},\mathbf{A}^{\mathrm{o}}\right)$ describes the correlation between options. The belief adjacency matrix $\mathbf{A}^{\mathrm{o}}=[a^{\mathrm{o}}_{jl}]\in\mathbb{R}^{N_{\mathrm{o}}\times N_{\mathrm{o}}}$ defines the strength of correlation between option $j$ and $l$.
4 Understanding Oversmoothing with Opinion Dynamics
In this section, we propose an analogy between GNNs and opinion dynamics models, and use our analogy to: (1) prove that recent continuous-depth GNN dynamics will exhibit oversmoothing for all inputs; and (2) design a novel continuous-depth GNN, behavior-inspired message passing (BIMP) GNN, using a nonlinear opinion dynamics inductive bias to circumvent oversmoothing and improve expressivity.
4.1 Oversmoothing is Opinion Consensus
Oversmoothing.
In the GNN literature, oversmoothing is loosely defined as the exponential convergence of node features with increasing network depth. A precise definition of oversmoothing is given in (Rusch et al., 2023).
The degree of oversmoothing in a network can be measured by the Dirichlet energy (Cai & Wang, 2020; Rusch et al., 2023) defined,
@@ 4.1 Oversmoothing is Opinion Consensus
where $\mathbf{x}^{l}_{i}$ denotes the feature associated with node $i$ at layer $l$. Oversmoothing is said to occur when the Dirichlet energy of layer $l$ tends to zero as $l$ tends to infinity, which is equivalent to the condition,
$\lim_{l\to\infty}\|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\|^{2}_{2}=0,\quad e_{ij}\in\mathcal{E}.$ (3)
Linear opinion dynamics.
Opinion dynamics models describe the evolution of agent opinions in a multi-agent system. The linear opinion dynamics model introduced in (DeGroot, 1974) describes the discrete-time evolution of agent opinions by
$\mathbf{x}_{i}(t+1)=\sum_{k=1}^{N_{\mathrm{a}}}a_{ik}\mathbf{x}_{k}(t),\quad\sum_{k=1}^{N_{\mathrm{a}}}a_{ik}=1,$ (4)
where $a_{ik}\geq 0$ denotes the influence of agent $i$ on agent $k$, and the total influence of any agent sums to one. The continuous-time analogue can be expressed,
$\dot{\mathbf{x}}_{i}(t)=-\mathbf{x}_{i}(t)+\sum_{k=1}^{N_{a}}a_{ik}\mathbf{x}_{k}(t).$ (5)
A limitation of linear opinion dynamics models, is that the opinions of agents tend toward consensus (i.e., agreement) for all initial conditions. Moreover, the consensus value is independent of the graph structure and linearly dependent on the initial conditions (Leonard et al., 2024).
Opinion consensus.
@@
An opinion dynamics model is said to reach consensus if and only if the opinions of all agents tend toward the same value as time tends to infinity, that is,
$\lim_{t\to\infty}\|\mathbf{x}_{i}(t)-\mathbf{x}_{j}(t)\|^{2}_{2}=0,\quad e_{ij}\in\mathcal{E}.$ (6)
GNN-OD Analogy
We construct an analogy between GNNs and opinion dynamics models to guide our analysis of existing continuous-time GNNs, and the development of our new continuous-time GNN, BIMP. Our analogy likens nodes at layer $l$ of a GNN with agents at time $t$ in a multi-agent system, node features with agent opinions, and oversmoothing with opinion consensus. The analogy is summarized in Table 1.
@@
Table 1: GNN-OD analogy. We construct an analogy between GNNs and opinion dynamics models. We relate the underlying graphical structures, and the criterion for oversmoothing and opinion concensus.

### v2

# Resolving Oversmoothing with Opinion Dissensus

@@ Abstract
While graph neural networks (GNNs) have allowed researchers to successfully apply neural networks to non-Euclidean domains, deep GNNs often exhibit lower predictive performance than their shallow counterparts. This phenomena has been attributed in part to oversmoothing, the tendency of node representations to become increasingly similar with network depth. In this paper we introduce an analogy between oversmoothing in GNNs and consensus (i.e., perfect agreement) in the opinion dynamics literature. We show that the message passing algorithms of several GNN models are equivalent to linear opinion dynamics models which have been shown to converge to consensus for all inputs regardless of the graph structure. This new perspective on oversmoothing motivates the use of nonlinear opinion dynamics as an inductive bias in GNN models. In our Behavior-Inspired Message Passing (BIMP) GNN, we leverage the nonlinear opinion dynamics model which is more general than the linear opinion dynamics model, and can be designed to converge to dissensus for general inputs. Through extensive experiments we show that BIMP resists oversmoothing beyond 100 time steps and consistently outperforms existing architectures even when those architectures are amended with oversmoothing mitigation techniques. We also show that BIMP has several desirable properties including well behaved gradients and adaptability to homophilic and heterophilic datasets.
@@ 1 Introduction
The underlying structure of a broad class of real-world systems is naturally captured by graphs. For example, in molecules, atoms can be represented by graph nodes and atomic bonds can be represented by edges [1]; in animal groups, individuals can be represented as nodes and their relative proximity can be represented by edges [2]; and in transportation networks, bus stops can be represented as nodes and public transit routes can be represented by edges [3]. Because of the broad applicability of graphs, their classification, regression, and generation are of strong interest across scientific communities. While MLPs can be adapted to operate on graph data, graph neural networks (GNNs) are specifically designed to respect the graph permutation symmetry (i.e., invariance to node relabeling) and can therefore learn generalizable graph representations, whereas MLPs are not.
GNNs have been successfully applied to tasks as diverse as recommendation [4], molecular prediction [5], protein design [6, 7], and computer chip design [8]. However, GNN representations often become increasingly similar with network depth; a phenomena known as oversmoothing [9, 10, 11]. A number of approaches have been proposed to address oversmoothing in GNNs, including the use of residual connections [9, 12, 13, 14], feature normalization [15, 16, 17], and alternative architectures [18, 19, 20]. While many alternative architectures work to incorporate higher-order features[21, 22, 23, 24], continuous-depth GNNs instead interpret conventional GNN architectures as discretizations of a continuous process [25, 26, 27]. This interpretation allows the application of techniques developed for modeling and analyzing dynamical systems [28, 29, 30, 31]. For example, GraphCON [32] avoids oversmoothing by enforcing stability conditions on hidden states of a coupled and damped oscillator system; and ACMP [33] avoids oversmoothing by introducing repulsive forces traditionally used to control clustering in particle systems. These are just a few among many that have shown incorporating inductive biases from physical processes can mitigate oversmoothing while maintaining expressivity [34].
In this paper, we propose a continuous-depth GNN inspired by behavioral interaction, instead of physical processes. First, we introduce an analogy likening node features in a GNN to opinions in an opinion dynamics model, feature aggregation to opinion exchange, and graph outputs to opinion outcomes (Section 3). Using this analogy, we show that oversmoothing will always occur in GNN models with layer-wise aggregation schemes that are equivalent to linear opinion dynamics (Section 4). With this new understanding of oversmoothing, we leverage the nonlinear opinion dynamics model [35, 36] to design a novel continuous-depth GNN that is provably robust to oversmoothing with desirable characteristics such as well behaved gradients and adaptability to heterophilic datasets (Section 5). Finally, we empirically validate our Behavior-Inspired Message Passing (BIMP) GNN on several datasets and against competitive baselines (Section 6).
@@ 2 Related Work
Oversmoothing in GNNs. Contrary to conventional feed forward networks [37, 38], deep discrete GNNs suffers degradation of performance from oversmoothing of node features [9, 10, 11]. Further analysis across wide classes of GNN models have resulted in ways to understand this phenomenon. Focusing on specific architectures, works have revealed that simple linear GNNs exhibits oversmoothing as the added layers exacerbate the denoising and mixing effect of the network [39]; GCNs [40] learns representations that attempts to counteract an inherently oversmoothing prone network structure [41]; and attention based networks such as GAT [42] oversmooths at an exponential rate due to the ergodicity of infinite matrix products [43]. More generally, oversmoothing have been characterized by energy minimization of gradient flows [44], representational rank collapse [45], and exceeding a theoretical limit of smoothing in mean aggregation [46].
Continuous-depth GNNs. Continuous-depth networks such as NeuralODE [47] define the network depth implicitly through the simulation of differential equations. GDE [25] leverages this notion to construct continuous-depth GNNs by propagating inputs through continuum of GNN layers governed by an underlying ODE. In order to better control and understand node dynamics, further work focuses on leveraging physical inductive biases such as heat diffusion [26], Beltrami flows [22], wave equations [27], coupled damped oscillators [32], energy source terms [48], Allen-Cahn reaction diffusion processes [33], blurring-sharpening forces [49], oscillator synchronization [50], and Ricci flow [51]. These embedded dynamics provide a systemic approach to counteract known limitations of discrete GNNs such as oversmoothing of deep networks and degradation of performance on heterophilic graphs [34].
@@
Opinion dynamics. Since opinions can be interpreted as latent preferences, modeling opinion dynamics provide insights into understanding and predicting agents’ behavior. Specifically, consensus dynamics [52] is commonly used in multi-agent settings such as coordinating multi-vehicle movements [53, 54], understanding network systems [55, 56], and learning on graphs [57]. However, the naive implementation of linear models for opinion formulation results in dynamics that only converge to consensus [58, 59]. This noted short comings of consensus dynamics models are resolved in nonlinear opinion dynamics [35]. The nonlinearity results in bifurcations [60] and hence opinions can evolve to dissensus rapidly even under weak input signals [36]. Nonlinear opinion dynamics have been shown to able to model a variety of systems such as group decision making [61, 62, 63]; multi-agent control [54, 64]; and relational inference [65]. Our proposed network seeks to integrate nonlinear opinion dynamics as the behavior-inspired inductive bias in a continuous-depth GNN.
3 Graph Neural Networks and Opinion Dynamics
In this section, we introduce our GNN-opinion dynamics (GNN-OD) analogy. We begin with a review GNNs and opinion dynamics, followed by a brief discussion of bifurcation theory in the context of the nonlinear opinion dynamics model [35]. We then develop an analogy likening GNNs to opinion dynamics models, and oversmoothing to opinion consensus.
3.1 Graph Neural Networks and Oversmoothing
Graph neural networks. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $n=|\mathcal{V}|$ nodes and $m=|\mathcal{E}|$ edges, where an edge $e_{ij}$ exists in $\mathcal{E}$ if the nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are connected in $\mathcal{G}$. Given an input graph, a graph neural network (GNN) $f:\mathcal{G}\rightarrow\mathcal{Y}$ returns a label (or label set) over edges, nodes, or the entire graph. Of the existing GNN algorithms, a large subset can be described in the message passing framework [66]. In this framework, layer-wise transformations are determined by learned message and update functions. The message function at layer $l$, $M^{l}$, and update function at layer $l$, $U^{l}$, are of the form,
$\mathbf{m}^{l+1}_{i}=\sum_{\mathbf{x}_{j}\in\mathcal{N}(\mathbf{x}_{i})}M^{l}(\mathbf{x}^{l}_{i},\mathbf{x}^{l}_{j},e_{ij})\quad\text{and}\quad\mathbf{x}^{l+1}_{i}=U^{l}(\mathbf{x}^{l}_{i},\mathbf{m}^{l+1}_{i}),$ (1)
where $\mathbf{x}^{l}_{i}$ denotes the representation of node $i$ at layer $l$.
Oversmoothing. In the GNN literature, oversmoothing is defined as the tendency for node features to become increasingly similar with increasing network depth [67]. This phenomena has been observed in discrete [39, 41, 43, 46] and continuous-depth [26, 27, 68] GNNs, and correlates with reduced predictive performance. We can measure the degree of oversmoothing in a GNN using the Dirichlet energy [67, 69]. The Dirichlet energy at layer $l$ is given by,
@@ 3.1 Graph Neural Networks and Oversmoothing
where $\mathbf{X}^{l}=[\mathbf{x}_{1}^{l},\cdots,\mathbf{x}_{n}^{l}]^{T}$. If the Dirichlet energy tends to zero as $l$ tends to infinity, that is,
$\lim_{l\to\infty}\|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\|^{2}_{2}=0\;\text{for all}\;e_{ij}\in\mathcal{E},$ (3)
the network is said to exhibit oversmoothing.
3.2 Opinion Dynamics and Opinion Consensus
Let $\mathcal{M}=(\mathcal{G}^{a},\mathcal{G}^{o})$ be a multi-agent system with a communication graph $\mathcal{G}^{\mathrm{a}}=(\mathcal{V}^{\mathrm{a}},\mathcal{E}^{\mathrm{a}},\mathbf{A}^{\mathrm{a}})$, and an option graph $\mathcal{G}^{\mathrm{o}}=\left(\mathcal{V}^{\mathrm{o}},\mathcal{E}^{\mathrm{o}},\mathbf{A}^{\mathrm{o}}\right)$. In this system, each of the $N_{\mathrm{a}}$ agents has a real-valued opinion on each of the $N_{\mathrm{o}}$ options. The adjacency matrix of the communication graph, $\mathbf{A}^{\mathrm{a}}=\left[a^{\mathrm{a}}_{ik}\right]\in\mathbb{R}^{N_{\mathrm{a}}\times N_{\mathrm{a}}}$, defines the communication strength between agents, and the adjacency matrix of the option graph, $\mathbf{A}^{\mathrm{o}}=[a^{\mathrm{o}}_{jl}]\in\mathbb{R}^{N_{\mathrm{o}}\times N_{\mathrm{o}}}$, defines how correlated different options are. An opinion dynamics model on $\mathcal{M}=(\mathcal{G}^{a},\mathcal{G}^{o})$ describes the evolution of agent opinions in time.
@@ 3.2 Opinion Dynamics and Opinion Consensus
Opinion consensus. Given a multi-agent system $\mathcal{M}=(\mathcal{G}^{a},\mathcal{G}^{o})$, a question of interest is whether the opinions of agents tend toward consensus (i.e., perfect agreement). The opinions of agents is said to reach consensus if and only if the opinions of all agents tend toward the same value as time tends to infinity [70], that is,
$\lim_{t\to\infty}\|\mathbf{x}_{i}(t)-\mathbf{x}_{j}(t)\|^{2}_{2}=0,\;\text{forall}\;e_{ij}\in\mathcal{E}.$ (4)
Linear opinion dynamics. The linear opinion dynamics model introduced in [70] describes the discrete-time evolution of agent opinions by,
$\mathbf{x}_{i}(t+1)=\sum_{k=1}^{N_{\mathrm{a}}}a^{\mathrm{a}}_{ik}\mathbf{x}_{k}(t),\quad\sum_{k=1}^{N_{\mathrm{a}}}a^{\mathrm{a}}_{ik}=1,$ (5)
where $a^{\mathrm{a}}_{ik}\geq 0$ can be interpreted as the influence of agent $x_{i}$ on agent $x_{k}$, and the total influence of any agent sums to one. In this model, the option graph can be understood as uncorrelated, $a^{\mathrm{o}}_{jl}=0$. The continuous-time analogue of Equation (5) is given by,
$\dot{\mathbf{x}}_{i}(t)=-d_{i}\mathbf{x}_{i}(t)+\sum_{k=1}^{N_{a}}a^{\mathrm{a}}_{ik}\mathbf{x}_{k}(t),$ (6)
where the total influence of agent $i$ is equal to $d_{i}$ [35].
In a linear opinion dynamics model, consensus is reached for all initial conditions, and the consensus value is independent of the graph structure and linearly dependent on the initial conditions [35]. The more general case of linear opinion dynamics with time-varying influence (i.e., $a^{\mathrm{a}}_{ik}(t)$ is time-dependent) can also be shown to converge to consensus [71, 72, 73, 74].
Figure 1: Nonlinear opinion dynamics and dissensus. (Left) The pitchfork bifurcation diagram illustrates a change in the number and stability of opinion states with the attention parameter $u$ (stable equilibria are illustrated with a solid line and unstable equilibria are illustrated with a dotted line). In the diagram, $z$ represents the weighted average of agent opinions, and $u^{*}$ represents the bifurcation point. When the input term $b=0$ we have the pitchfork bifurcation (top-left), and when $b>0$ we have its unfolding (bottom-left). (Right) The time evolution of agent opinions under the nonlinear opinion dynamics model depends on the initial weighted average of agent opinions $z$, the attention parameter $u$, and the weighted average of agent inputs $b$. Each subfigure corresponds to an initial condition on the left. In all cases, the initial $z$ is the same. (a) ($u<u^{*}$, $b=0$). Agent opinions converge to a neutral consensus (i.e., perfect agreement) which is equivalent to oversmoothing. (b) ($u>u^{*}$, $b=0$). Agent opinions converge to dissensus with low variance, $z$ is positive. (c) ($u=u^{*}$, $b=0$). Agent opinions converge to a neutral consensus. (d) ($u=u^{*}$, $b>0$). Agent opinions converge to dissensus with high variance, $z$ is positive.
Nonlinear opinion dynamics. The nonlinear opinion dynamics model introduced in [35, 36, 75] describes the continuous-time evolution of agent $i$’s opinion about option $j$ by,
$\dot{x}_{ij}=-d_{ij}x_{ij}+S\Bigg{(}u_{i}\Bigg{(}\alpha_{ij}x_{ij}+\sum^{N_{\mathrm{a}}}_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}a^{\mathrm{a}}_{ik}x_{kj}+\sum^{N_{\mathrm{o}}}_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}a^{\mathrm{o}}_{jl}x_{il}+\sum^{N_{\mathrm{a}}}_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}\sum^{N_{\mathrm{o}}}_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{jl}x_{kj}\Bigg{)}\Bigg{)}+b_{ij},$ (7)
where $d_{ij}\geq 0$, $u_{i}>0$, and $\alpha_{ij}\geq 0$ are parameters intrinsic to the agent, and $a^{\mathrm{a}}_{jk}$, $a^{\mathrm{o}}_{jl}$, and $b_{ij}$ are parameters extrinsic to the agent. The intrinsic parameter $d_{ij}$, the damping parameter, describes how resistant agent $i$ is to forming an opinion about option $j$; $u_{i}$, the attention parameter, represents the attentiveness of agent $i$ to the opinions of other agents; and $\alpha_{ij}$, the self-reinforcement parameter, defines how confident agent $i$ is in its opinion on option $j$. The extrinsic parameter $b_{ij}$, the input parameter, represents the impact of the environment on the agent $i$’s opinion about option $j$, and the saturating function $\mathcal{S}$ is selected so that $S\left(0\right)=0$, $S^{\prime}\left(0\right)=1$, and $S^{\prime\prime\prime}\left(0\right)\neq 0$.
By modeling opinion formation as a nonlinear process, the nonlinear opinion dynamics model can capture opinion consensus and dissensus. The nonlinearity induces a bifurcation where the number and/or stability of equilibrium solutions changes (see Figure 1 top-left). Consensus results when all agents select exactly the same equilibrium value, dissensus results otherwise. A switch from consensus to dissensus can result from a change in the attention parameter $u$, this can be seen by comparing the dynamics at point (a) and (b) in Figure 1 (left), or from a change in the input parameter $b$, this can be seen by comparing the dynamics at point (c) and (d) in Figure 1 (left).
@@
Table 1: GNN-OD Analogy. We describe our analogy between GNNs and opinion dynamics. Our analogy relates the graph structures in GNNs and opinion dynamics, and the notions of oversmoothing and opinion consensus.